# EDEN: Exploring Disks Embedded in N-body simulations of Milky-Way-mass halos from Symphony

### v1

@@ 4.2 Subhalo peak mass functions
We also compare the median SPMF of the nine halos in the EDEN $M_{\ast,\rm Disk}\times 2.5$ subsample with their Symphony DMO counterparts. Their DMO counterparts, labeled as ‘Symphony $M_{\ast,\rm Disk}\times 2.5$’, show a similar median SPMF to the full Symphony suite, suggesting that this nine-halo subsample is an unbiased representation of the 45 Symphony halos. However, the SPMF of the EDEN $M_{\ast,\rm Disk}\times 2.5$ halos are more suppressed than the EDEN fiducial median SPMF at all resolved subhalo masses for both $r<100$ kpc and $r<R_{\rm vir,host}$, which is another manifestation of subhalo abundance suppression being sensitive to the disk-to-halo mass ratio (Fig. 5).
Figure 7: Subhalo radial cumulative distributions for resolved subhalos with $M_{\rm peak}\geqslant 300m_{\rm DM}$ ($1.2\times 10^{8}\mathrm{M_{\astrosun}}$). The left panels show the subhalo abundance versus the $z=0$ distance ($r$) to the MW host center, and the right panels show the radial cumulative function of their orbit pericenter distances ($d_{\rm peri}$). Solid curves show average subhalo radial distributions for the Symphony (blue) and EDEN (red); shaded regions with the same color denote the corresponding $68\%$ host-to-host scatter. The dashed red curve shows the average subhalo radial distribution for the nine halos in EDEN $M_{\ast,\rm Disk}\times 2.5$, and the blue dashed curve shows the corresponding DMO subsample in Symphony. The lower side panels show the CDF ratio of each curve relative to Symphony. The triangles in the right panel denote the scale radii of the ‘stellar’ (S, 2.5 kpc) and ‘gaseous’ (G, 7 kpc) disks of the embedded potential. Subhalo abundance suppression becomes weaker for subhalos penetrating into the inner parts of the disk potential. We caution that these radial distributions may change significantly due to numerical resolution (Mansfield et al., 2023).
@@ 4.3 Subhalo radial distribution
Figure 8: Subhalo counts ratio as a function of their first-infall lookback-time into the MW host for resolved subhalos with $M_{\rm peak}\geqslant 300\,m_{\rm DM}$ ($1.2\times 10^{8}\mathrm{M_{\astrosun}}$). The left panel shows subhalos within 100 kpc, and the right panel shows subhalos within $R_{\rm vir,host}$. The orange circles denote the EDEN fiducial halos, and the turquoise diamonds denote the nine halos in EDEN $M_{\ast,\rm Disk}\times 2.5$. The shaded bands are Poisson errors on the subhalo count ratios. The subhalos removed from our sample by the inclusion of a disk are preferentially those with earlier infall times.
To summarize, halos with disks have fewer subhalos that fall on orbits with pericenter distances $d_{\rm peri}\lesssim 100$ kpc. Increasing the disk mass increases subhalo abundance suppression strength, especially at $d_{\rm peri}\lesssim 50$ kpc. We caution that these radial trends in subhalo abundances should be interpreted as qualitative instead of quantitative as subhalo radial distributions can change significantly (Mansfield et al., 2023).
@@ 4.4 Subhalo infall times
As we have shown in the previous sections, subhalo abundance suppression in the presence of a galactic disk is sensitive to the disk-to-halo mass ratio, the subhalo peak mass, and subhalo orbit pericenter distances. As found in Phat ELVIS (Kelley et al., 2019), subhalos that fell in earlier to their hosts are more preferentially stripped than later-infalling subhalos (see also Fig. 1 in Jiang & van den Bosch 2017). The physical intuition behind this effect is that earlier-infalling subhalos fell in when their host halos were smaller and have systematically smaller pericenters (Wetzel & White, 2010), leading to stronger tidal stripping from the disk that enhance their mass loss. In the following, we investigate if this effect is generic in EDEN as Phat ELVIS used a MW-like heavy disk mass (Fig. 1).
In Fig. 8, we show the subhalo counts ratio in EDEN versus Symphony as a function of the lookback time of the first-infall into their MW-mass hosts ($r<100$ kpc in the left panel and $r<R_{\rm vir,host}$ in the right panel). In both radial ranges, subhalo abundance suppression mainly happens for early-infalling subhalos with $t_{\rm first,infall}>8$ Gyrs, consistent with the findings in Phat ELVIS (Kelley et al., 2019). There are about $\sim 10\%$ of subhalos with $r<100$ kpc and fell in from $4\sim 8$ Gyrs ago that are lost from disk tidal stripping. However, this suppression is not seen considering all subhalos within $R_{\rm vir,host}$ that fell in $4\sim 8$ Gyrs ago. This indicates that the additional subhalos brought in by adiabatic contraction (Fig. 5) that have not yet entered the central 100 kpc of the host have not been severely stripped and compensates for the central subhalo abundance suppression (Fig. 7). The nine halos in EDEN $M_{\ast,\rm Disk}\times 2.5$ further enhance the infall time distribution patterns of EDEN fiducial, with most of the additional suppression of subhalo abundance occurring with $t_{\rm first,infall}>8$ Gyrs due to the increase in disk-to-halo mass ratio. These findings indicate that the MW disk predominantly loses early-infalling subhalos due to disk tidal stripping, and increasing the disk mass further enhances this effect.
@@ 5.1 A user’s guide to interpreting the scope of theoretical subhalo disruption/mass loss studies
Each selection choice has its own considerations. First, they target different classes of observations: present-day measures of mass are good tracers of observations that directly probe present-day subhalo mass (e.g., stream gaps, strong lensing, rotation curves, etc.; Bechtol et al. 2022), while pre-infall proxies tend to be better tracers of satellite galaxies selected by stellar mass (e.g., Reddick et al., 2013). Second, population demographics can be very different for samples selected by pre-infall mass proxies than present-day mass proxies. This is particularly true for the radial distribution of subhalos, which is more concentrated in subhalos selected by peak mass than by present mass (e.g., Nagai & Kravtsov, 2005; Han et al., 2016). Third, numerical concerns are different for different mass proxies (Mansfield et al., 2023): i) pre-infall selections have more stringent resolution requirements than present-day mass selections and may need to contend with aphysical fluctuations in mass as subhalos disrupt (see also van den Bosch 2017); ii) present-day mass measurements need to contend with the fact that different subhalo finders have different ways of separating the mass of a subhalo from its host and different schemes can bias masses by tens-of-percent; iii) maintaining unbiased $v_{\rm max}$ values in subhalo velocity profiles for subhalos undergoing mass loss requires almost an order of magnitude more resolution than maintaining unbiased $M_{\rm vir}$ values.
@@
Resolution: A poorly resolved subhalo may not give accurate predictions for the $\Lambda$CDM model. As subhalos lose mass, they experience more classical two-body scattering (e.g., Power et al., 2003; Ludlow et al., 2019); they may further experience decreased durability due to their simulation’s force softening (van den Bosch & Ogiya, 2018; Mansfield & Avestruz, 2021), and can eventually encounter runaway mass loss due to discreteness effects (van den Bosch & Ogiya, 2018). In a well-calibrated simulation, all three effects affect subhalos at similar resolution levels (Mansfield et al., 2023), but this may not be true in a simulation with excessively large or small force softening scales. Certain subhalo population statistics are strongly dependent on resolution, such as the abundance or radial distribution of subhalos selected at a fixed $M_{\rm peak}$.
@@
Subhalo finder: Subhalo finders are imperfect tools and are often the dominant source of numerical biases in subhalo analysis. For example, the Rockstar subhalo finder (Behroozi et al., 2013b) is one of the most widely used subhalo finders and generally performs at least as well as other major subhalo finders in comparative tests between tools (Knebe et al., 2011; Onions et al., 2012, 2013; Srisawat et al., 2013; Avila et al., 2014; Behroozi et al., 2014; Elahi et al., 2019). However, Rockstar demonstrably loses track of even high-resolution subhalos after relatively modest amounts of mass loss (Griffen et al., 2016; Mansfield et al., 2023) and falsely converges (Mansfield et al., 2023), meaning that even as resolution is increased and the simulation’s subhalos are able to survive for longer, Rockstar will continue losing track of subhalos after they have lost the same amount of mass. The strengths and drawbacks of every halo finder are different and it is important to understand what those weaknesses are in depth, even for qualitative studies. For some subhalo finders, these limitations have never been formally studied. We also recommend that one should generally be skeptical of reliability arguments based on subhalo finder comparisons or internal convergence tests, given that such tests had been performed extensively on the Rockstar subhalo finder without identifying its aforementioned issues.
@@
Galaxy Mass Loss/Orphan Model To first order, particles can be stripped from a subhalo when they orbit close to that subhalo’s tidal radius, and a subhalo’s tidal radius gradually shrinks as it orbits its host. Because galaxies are small compared to their halos, low-mass satellite galaxies generally stay intact through their initial mass loss and begin to rapidly lose mass after passing a critical threshold in $M_{\rm vir}/M_{\rm peak}$ (Smith et al., 2013). Whether or not this rapid mass loss is qualitatively consistent with “disruption” depends on the stellar density profile of the satellite (Errani et al., 2023b). For subhalo finders that lose track of subhalos after this threshold is reached (e.g. Symfind, Mansfield et al., 2023), a galaxy mass loss model is needed, lest the finder over predict the abundance of satellite galaxies at a fixed stellar mass. For subhalo finders that lose track of subhalos prior to this threshold (e.g. Rockstar Mansfield et al., 2023) require both a galaxy disruption model and an “orphan” model, i.e., a model that creates post hoc galaxy tracers and evolves them forward based on heuristics. Several points to consider when evaluating these models are (a) that mass loss models calibrated on hydrodynamic simulations likely vastly overestimate mass loss rates due to numerical issues (e.g., Ludlow et al., 2019), (b) that dynamical friction is quite weak in low mass subhalos (van den Bosch et al., 2016) meaning that it is unlikely to be their primary disruption mechanism despite it being a popular prescription, and (c) that the popular choice of generating orphan galaxies during a subhalo’s last snapshot is questionable given that the subhalo recovered during this snapshot is often extremely unreliable.
@@
The interplay between the numerics of our simulations and our chosen subhalo finder, Symfind, have been extensively explored in Mansfield et al. (2023). Symfind is able to track subhalos well past the point where their galaxies have started to rapidly lose mass and will not be a limiting factor in analysis. However, some subhalos in the $n_{\rm peak}\lesssim 5^{3}$ range will experience some degree of accelerated mass loss due to simulation resolution prior to the point when stellar mass loss should begin, although we do not currently have a model that can make quantitative estimates for how many subhalos will be missing in this range. The use of Symfind significantly increases the number of subhalos that we find relative to Rockstar (see Appendix D), and the choice between subhalo finders is comparable in effect to the amount of suppression introduced by the disk itself. Much of this difference is due to Rockstar losing track of subhalos early, but some unknown fraction of the difference may come from the fact that Symfind is holding onto subhalos so long that their satellite galaxies would have lost enough stellar mass to move outside the equivalent stellar-mass-selected sample. Future work on precise stellar mass loss models will help resolve this issue.
@@ 5.2 Comparisons with previous embedded disk simulations
We conjecture two possible factors that can potentially lead to the stronger subhalo abundance suppression seen in Garrison-Kimmel et al. (2017). One possibility is that the subhalo orbits of these two halos had smaller pericenters and were accreted more radially due to the large-scale environment of the host (e.g., accreting subhalos collectively from filaments). Another possibility is that the fragility of AHF halo finder used by Garrison-Kimmel et al. (2017) has even worse subhalo tracking issues than Rockstar (Mansfield et al., 2023), leading to more heavily stripped subhalos being artificially lost track of in the presence of a disk potential. A similar effect could be applicable to AHF (Knollmann & Knebe, 2009), the halo finder used in Garrison-Kimmel et al. (2017), such that it numerically enhances the suppression of subhalo abundance. Since only two halos were picked from FIRE-2 and they already happen to yield drastically lower $N_{\rm Disk}/N_{\rm DMO}$, the average mass-loss for subhalos in FIRE-2 is likely to be statistically stronger than that of EDEN (combining physical and numerical effects). We defer a more careful analysis of this issue to future work.
Regarding subhalo radial distributions, EDEN qualitatively agrees with Garrison-Kimmel et al. (2017) and Kelley et al. (2019). However, there are a few quantitative differences we wish to highlight. Comparing EDEN with Garrison-Kimmel et al. (2017), the latter (see their Fig. 3) has a similar 3D radial distribution to EDEN $M_{\ast,\rm Disk}\times 2.5$ (Fig. 7 left panel) and loses $>2/3$ subhalos within 30 kpc due to the added disk. The differences in 3D subhalo distances mainly come from the cumulative subhalo counts within 100 kpc being $\sim 15\%$ lower than EDEN $M_{\ast,\rm Disk}\times 2.5$ following discussion above. Comparing Phat ELVIS to EDEN, both Phat ELVIS and EDEN $M_{\ast,\rm Disk}\times 2.5$ do not have subhalos within 20 kpc of their hosts with the disk added with the suppression ratio decreasing at increasing distances (see their Fig. 3).
@@
Given that these two previous simulations reach similar pericenter distributions with different halo finders, it is most likely that the higher numerical resolution adopted in them leads to both having way more subhalos in their DMO runs with $d_{\rm peri}<20$ kpc compared to SymphonyMilkyWay (DMO). The higher number density in their DMO runs sets a higher $N_{\rm DMO}$, leading to a much lower $N_{\rm Disk}/N_{\rm DMO}$ compared to EDEN EDEN. As mentioned in Section 4.3, the subhalo radial distributions are sensitive to numerical resolution, and perhaps the pericenter distances are particularly more sensitive since it probes the most tidally violent part of each subhalo’s orbit that requires the most resolving power to numerically converge (Peñarrubia et al., 2010; van den Bosch & Ogiya, 2018; van den Bosch et al., 2018; Errani & Peñarrubia, 2020; Errani & Navarro, 2021; Mansfield et al., 2023). Please refer to Appendix D for a discussion of resolution effects.
Finally, EDEN results are also largely consistent with Green et al. (2022) in which various properties of the disk impact on NFW-profile (cuspy) subhalos are explored using the SatGen semi-analytic model (Jiang et al., 2021). They found that the disk mass is the most dominant factor in determining subhalo abundance suppression intensity over other factors such as disk scale radius and axis ratio. Within their framework, they find that heavier disks cause stronger subhalo abundance suppression, with $M_{\ast,\rm Disk}/M_{\rm vir,host}=0.05$ disks causing $\sim 30\%$ subhalos lost within $50$ kpc and $\sim 10\%$ subhalos lost within $R_{\rm vir,host}$, qualitatively agreeing with the EDEN results (Fig. 5). They also find that the radial distribution of subhalos are more suppressed in the inner regions within $\sim 100$ kpc with increasing disk mass, similar to EDEN (Fig. 7). These findings further corroborate the key role disk-to-halo mass ratio plays in modeling the tidal stripping and mass loss of subhalos due to the central galaxy.
We note that in the case of cored-subhalos due to internal stellar feedback as in hydrodynamic simulations (e.g., Garrison-Kimmel et al., 2017; Samuel et al., 2020) or emulated semi-analytic models (e.g., Jiang et al., 2021), subhalo abundance suppression could be more intense due to the less-concentrated nature of the subhalo itself as compared to the cuspy DMO subhalos. This effect is best exemplified in the two halos from Garrison-Kimmel et al. (2017) where the subhalo abundance within 100 kpc in their hydrodynamic runs are always lower than their embedded disk runs. Therefore, more subhalo mass loss and abundance suppression could occur in reality than the EDEN predictions in this work.

### v3

@@ 4.2 Subhalo peak mass functions
We also compare the median SPMF of the nine halos in the EDEN $M_{\ast,\rm Disk}\times 2.5$ subsample with their Symphony DMO counterparts. Their DMO counterparts, labeled as ‘Symphony $M_{\ast,\rm Disk}\times 2.5$’, show a similar median SPMF to the full Symphony suite, suggesting that this nine-halo subsample is an unbiased representation of the 45 Symphony halos. However, the SPMF of the EDEN $M_{\ast,\rm Disk}\times 2.5$ halos is more suppressed than the EDEN fiducial median SPMF at all resolved subhalo masses for both $r<100$ kpc and $r<R_{\rm vir,host}$, which is another manifestation of subhalo abundance suppression being sensitive to the disk-to-halo mass ratio (Fig. 5).
Figure 7: Subhalo radial cumulative distributions for resolved subhalos with $M_{\rm peak}\geqslant 300m_{\rm DM}$ ($1.2\times 10^{8}\mathrm{M_{\astrosun}}$). The left panels show the subhalo abundance versus the $z=0$ distance ($r$) to the MW host center, and the right panels show the radial cumulative function of their orbit pericenter distances ($d_{\rm peri}$). Solid curves show average subhalo radial distributions for the Symphony (blue) and EDEN (red); shaded regions with the same color denote the corresponding $68\%$ host-to-host scatter. The dashed red curve shows the average subhalo radial distribution for the nine halos in EDEN $M_{\ast,\rm Disk}\times 2.5$, and the blue dashed curve shows the corresponding DMO subsample in Symphony. The lower side panels show the CDF ratio of each curve relative to Symphony. The triangles in the right panel denote the scale radii of the ‘stellar’ (S, 2.5 kpc) and ‘gaseous’ (G, 7 kpc) disks of the embedded potential. Subhalo abundance suppression becomes weaker for subhalos penetrating into the inner parts of the disk potential. We caution that these radial distributions may change significantly due to numerical resolution (Mansfield et al., 2024).
@@ 4.3 Subhalo radial distribution
Figure 8: Subhalo count ratio as a function of their first-infall lookback-time into the MW host for resolved subhalos with $M_{\rm peak}\geqslant 300\,m_{\rm DM}$ ($1.2\times 10^{8}\mathrm{M_{\astrosun}}$). The left panel shows subhalos within 100 kpc, and the right panel shows subhalos within $R_{\rm vir,host}$. The orange circles denote the EDEN fiducial halos, and the turquoise diamonds denote the nine halos in EDEN $M_{\ast,\rm Disk}\times 2.5$. The shaded bands are Poisson errors on the subhalo count ratios. The subhalos removed from our sample by the inclusion of a disk are preferentially those with earlier infall times.
To summarize, halos with disks have fewer subhalos that fall on orbits with pericenter distances $d_{\rm peri}\lesssim 100$ kpc. Increasing the disk mass increases subhalo abundance suppression strength, especially at $d_{\rm peri}\lesssim 50$ kpc. We caution that these radial trends in subhalo abundances should be interpreted as qualitative instead of quantitative as subhalo radial distributions can change significantly (Mansfield et al., 2024).
@@ 4.4 Subhalo infall times
As we have shown in the previous sections, subhalo abundance suppression in the presence of a galactic disk is sensitive to the disk-to-halo mass ratio, the subhalo peak mass, and subhalo orbit pericenter distances. As found in Phat ELVIS (Kelley et al., 2019), subhalos that fell in earlier to their hosts are more preferentially stripped than later-infalling subhalos (see also Fig. 1 in Jiang & van den Bosch 2017). The physical intuition behind this effect is that earlier-infalling subhalos fell in when their host halos were smaller and have systematically smaller pericenters (Wetzel & White, 2010), leading to stronger tidal stripping from the disk that enhances their mass loss. In the following, we investigate if this effect is generic in EDEN as Phat ELVIS used a MW-like heavy disk mass (Fig. 1).
In Fig. 8, we show the subhalo count ratio in EDEN versus Symphony as a function of the lookback time of the first-infall into their MW-mass hosts ($r<100$ kpc in the left panel and $r<R_{\rm vir,host}$ in the right panel). In both radial ranges, subhalo abundance suppression mainly happens for early-infalling subhalos with $t_{\rm first,infall}>8$ Gyrs, consistent with the findings in Phat ELVIS (Kelley et al., 2019). There are about $\sim 10\%$ of subhalos with $r<100$ kpc and fell in from $4\sim 8$ Gyrs ago that are lost from disk tidal stripping. However, this suppression effect is washed out if we consider all subhalos within $R_{\rm vir,host}$ that fell in $4\sim 8$ Gyrs ago. The nine halos in EDEN $M_{\ast,\rm Disk}\times 2.5$ further enhance the infall time distribution patterns of EDEN fiducial, with most of the additional suppression of subhalo abundance occurring with $t_{\rm first,infall}>8$ Gyrs due to the increase in disk-to-halo mass ratio. These findings indicate that the MW disk predominantly loses early-infalling subhalos due to disk tidal stripping, and increasing the disk mass further enhances this effect.
@@ 5.1 A user’s guide to interpreting the scope of theoretical subhalo disruption/mass loss studies
Each selection choice has its own considerations. First, they target different classes of observations: present-day measures of mass are good tracers of observations that directly probe present-day subhalo mass (e.g., stream gaps, strong lensing, rotation curves, etc.; Bechtol et al. 2022), while pre-infall proxies tend to be better tracers of satellite galaxies selected by stellar mass (e.g., Reddick et al., 2013). Second, population demographics can be very different for samples selected by pre-infall mass proxies than present-day mass proxies. This is particularly true for the radial distribution of subhalos, which is more concentrated in subhalos selected by peak mass than by present mass (e.g., Nagai & Kravtsov, 2005; Han et al., 2016). Third, numerical concerns are different for different mass proxies (Mansfield et al., 2024): i) pre-infall selections have more stringent resolution requirements than present-day mass selections and may need to contend with aphysical fluctuations in mass as subhalos disrupt (see also van den Bosch 2017); ii) present-day mass measurements need to contend with the fact that different subhalo finders have different ways of separating the mass of a subhalo from its host and different schemes can bias masses by tens-of-percent; iii) maintaining unbiased $v_{\rm max}$ values in subhalo velocity profiles for subhalos undergoing mass loss requires almost an order of magnitude more resolution than maintaining unbiased $M_{\rm vir}$ values.
@@
Resolution: A poorly resolved subhalo may not give accurate predictions for the $\Lambda$CDM model. As subhalos lose mass, they experience more classical two-body scattering (e.g., Power et al., 2003; Ludlow et al., 2019b); they may further experience decreased durability due to their simulation’s force softening (van den Bosch & Ogiya, 2018; Mansfield & Avestruz, 2021), and can eventually encounter runaway mass loss due to discreteness effects (van den Bosch & Ogiya, 2018). In a well-calibrated simulation, all three effects affect subhalos at similar resolution levels (Mansfield et al., 2024), but this may not be true in a simulation with excessively large or small force softening scales. Certain subhalo population statistics are strongly dependent on resolution, such as the abundance or radial distribution of subhalos selected at a fixed $M_{\rm peak}$.
@@
Subhalo finder: Subhalo finders are imperfect tools and are often the dominant source of numerical biases in subhalo analysis. For example, the Rockstar subhalo finder (Behroozi et al., 2013b) is one of the most widely used subhalo finders and generally performs at least as well as other major subhalo finders in comparative tests between tools (Knebe et al., 2011; Onions et al., 2012, 2013; Srisawat et al., 2013; Avila et al., 2014; Behroozi et al., 2014; Elahi et al., 2019). However, Rockstar demonstrably loses track of even high-resolution subhalos after relatively modest amounts of mass loss (Griffen et al., 2016; Mansfield et al., 2024) and falsely converges (Mansfield et al., 2024), meaning that even as resolution is increased and the simulation’s subhalos are able to survive for longer, Rockstar will continue losing track of subhalos after they have lost the same amount of mass. The strengths and drawbacks of every halo finder are different, and it is important to understand what those weaknesses are in depth, even for qualitative studies. For some subhalo finders, these limitations have never been formally studied. We also recommend that one should generally be skeptical of reliability arguments based on subhalo finder comparisons or internal convergence tests, given that such tests have been performed extensively on the Rockstar subhalo finder without identifying its aforementioned issues.
@@
Galaxy Mass Loss/Orphan Model To first order, particles can be stripped from a subhalo when they orbit close to that subhalo’s tidal radius, and a subhalo’s tidal radius gradually shrinks as it orbits its host. Because galaxies are small compared to their halos, low-mass satellite galaxies generally stay intact through their initial mass loss and begin to rapidly lose mass after passing a critical threshold in $M_{\rm vir}/M_{\rm peak}$ (Smith et al., 2013). Whether or not this rapid mass loss is qualitatively consistent with “disruption” depends on the stellar density profile of the satellite (Errani et al., 2023). For subhalo finders that lose track of subhalos after this threshold is reached (e.g. Symfind, Mansfield et al., 2024), a galaxy mass loss model is needed, lest the finder overpredict the abundance of satellite galaxies at a fixed stellar mass. For subhalo finders that lose track of subhalos prior to this threshold (e.g. Rockstar Mansfield et al., 2024) require both a galaxy disruption model and an “orphan” model, i.e., a model that creates post hoc galaxy tracers and evolves them forward based on heuristics. Several points to consider when evaluating these models are (a) that mass loss models calibrated on hydrodynamic simulations likely vastly overestimate mass loss rates due to numerical issues (e.g., Ludlow et al., 2019b), (b) that dynamical friction is quite weak in low-mass subhalos (van den Bosch et al., 2016) meaning that it is unlikely to be their primary disruption mechanism despite it being a popular prescription, and (c) that the popular choice of generating orphan galaxies during a subhalo’s last snapshot is questionable given that the subhalo recovered during this snapshot is often extremely unreliable.
@@
The interplay between the numerics of our simulations and our chosen subhalo finder, Symfind, has been extensively explored in Mansfield et al. (2024). Symfind is able to track subhalos well past the point where their galaxies have started to rapidly lose mass and will not be a limiting factor in analysis. However, some subhalos in the $n_{\rm peak}\lesssim 5^{3}$ range will experience some degree of accelerated mass loss due to simulation resolution prior to the point when stellar mass loss should begin, although we do not currently have a model that can make quantitative estimates for how many subhalos will be missing in this range. The use of Symfind significantly increases the number of subhalos that we find relative to Rockstar (see Appendix D), and the choice between subhalo finders is comparable in effect to the amount of suppression introduced by the disk itself. Much of this difference is due to Rockstar losing track of subhalos early, but some unknown fraction of the difference may come from the fact that Symfind is holding onto subhalos so long that their satellite galaxies would have lost enough stellar mass to move outside the equivalent stellar-mass-selected sample. Future work on precise stellar mass loss models will help resolve this issue.
@@ 5.2 Comparisons with previous embedded disk simulations
We conjecture two possible factors that can potentially lead to the stronger subhalo abundance suppression seen in Garrison-Kimmel et al. (2017). One possibility is that the subhalo orbits of these two halos had smaller pericenters and were accreted more radially due to the large-scale environment of the host (e.g., accreting subhalos collectively from filaments). Another possibility is that the fragility of AHF halo finder used by Garrison-Kimmel et al. (2017) has even worse subhalo tracking issues than Rockstar (Mansfield et al., 2024), leading to more heavily stripped subhalos being artificially lost track of in the presence of a disk potential. A similar effect could be applicable to AHF (Knollmann & Knebe, 2009), the halo finder used in Garrison-Kimmel et al. (2017), such that it numerically enhances the suppression of subhalo abundance. Since only two halos were picked from FIRE-2 and they already happen to yield drastically lower $N_{\rm Disk}/N_{\rm DMO}$, the average mass loss for subhalos in FIRE-2 is likely to be statistically stronger than that of EDEN (combining physical and numerical effects). We defer a more careful analysis of this issue to future work.
Regarding subhalo radial distributions, EDEN qualitatively agrees with Garrison-Kimmel et al. (2017) and Kelley et al. (2019). However, there are a few quantitative differences we wish to highlight. Comparing EDEN with Garrison-Kimmel et al. (2017), the latter (see their Fig. 3) has a similar 3D radial distribution to EDEN $M_{\ast,\rm Disk}\times 2.5$ (Fig. 7 left panel) and loses $>2/3$ subhalos within 30 kpc due to the added disk. The differences in 3D subhalo distances mainly come from the cumulative subhalo counts within 100 kpc being $\sim 15\%$ lower than EDEN $M_{\ast,\rm Disk}\times 2.5$ following the discussion above. Comparing Phat ELVIS to EDEN, both Phat ELVIS and EDEN $M_{\ast,\rm Disk}\times 2.5$ do not have subhalos within 20 kpc of their hosts, with the suppression ratio decreasing at increasing distances (see their Fig. 3).
@@
Given that these two previous simulations reach similar pericenter distributions with different halo finders, it is most likely that the higher numerical resolution adopted in them leads to both having way more subhalos in their DMO runs with $d_{\rm peri}<20$ kpc compared to SymphonyMilkyWay (DMO). The higher number density in their DMO runs sets a higher $N_{\rm DMO}$, leading to a much lower $N_{\rm Disk}/N_{\rm DMO}$ compared to EDEN EDEN. As mentioned in Section 4.3, the subhalo radial distributions are sensitive to numerical resolution, and perhaps the pericenter distances are particularly more sensitive since they probe the most tidally violent part of each subhalo’s orbit that requires the most resolving power to numerically converge (Peñarrubia et al., 2010; van den Bosch & Ogiya, 2018; van den Bosch et al., 2018; Errani & Peñarrubia, 2020; Errani & Navarro, 2021; Mansfield et al., 2024). Please refer to Appendix D for a discussion of resolution effects.
Finally, EDEN results are also largely consistent with Green et al. (2022) in which various properties of the disk impact on NFW-profile (cuspy) subhalos are explored using the SatGen semi-analytic model (Jiang et al., 2021). They found that the disk mass is the most dominant factor in determining subhalo abundance suppression intensity over other factors such as disk scale radius and axis ratio. Within their framework, they find that heavier disks cause stronger subhalo abundance suppression, with $M_{\ast,\rm Disk}/M_{\rm vir,host}=0.05$ disks causing $\sim 30\%$ subhalos lost within $50$ kpc and $\sim 10\%$ subhalos lost within $R_{\rm vir,host}$, qualitatively agreeing with the EDEN results (Fig. 5). They also find that the radial distribution of subhalos is more suppressed in the inner regions within $\sim 100$ kpc with increasing disk mass, similar to EDEN (Fig. 7). These findings further corroborate the key role disk-to-halo mass ratio plays in modeling the tidal stripping and mass loss of subhalos due to the central galaxy.
We note that in the case of cored-subhalos due to internal stellar feedback as in hydrodynamic simulations (e.g., Garrison-Kimmel et al., 2017; Samuel et al., 2020) or emulated semi-analytic models (e.g., Jiang et al., 2021), subhalo abundance suppression could be more intense due to the less-concentrated nature of the subhalo itself as compared to the cuspy DMO subhalos. This effect is best exemplified in the two halos from Garrison-Kimmel et al. (2017) where the subhalo abundance within 100 kpc in their hydrodynamic runs is always lower than their embedded disk runs. Therefore, more subhalo mass loss and abundance suppression could occur in reality than the EDEN predictions in this work.